# Uncertainty Matters in Dynamic Gaussian Splatting for Monocular 4D Reconstruction

### v1

@@ Abstract
Reconstructing dynamic 3D scenes from monocular input is fundamentally under-constrained, with ambiguities arising from occlusion and extreme novel views. While dynamic Gaussian Splatting offers an efficient representation, vanilla models optimize all Gaussian primitives uniformly, ignoring whether they are well or poorly observed. This limitation leads to motion drifts under occlusion and degraded synthesis when extrapolating to unseen views. We argue that uncertainty matters: Gaussians with recurring observations across views and time act as reliable anchors to guide motion, whereas those with limited visibility are treated as less reliable. To this end, we introduce USplat4D, a novel Uncertainty-aware dynamic Gaussian Splatting framework that propagates reliable motion cues to enhance 4D reconstruction. Our key insight is to estimate time-varying per-Gaussian uncertainty and leverages it to construct a spatio-temporal graph for uncertainty-aware optimization. Experiments on diverse real and synthetic datasets show that explicitly modeling uncertainty consistently improves dynamic Gaussian Splatting models, yielding more stable geometry under occlusion and high-quality synthesis at extreme viewpoints. Project page: https://tamu-visual-ai.github.io/usplat4d/.
Figure 1: We show a challenging case from the DyCheck dataset (Gao et al., 2022), where a person casually rotates a backpack while being captured by a moving monocular camera. The goal is to reconstruct the dynamic object for arbitrary viewpoints and timestamps. The state-of-the-art dynamic Gaussian Splatting methods, e.g., Shape-of-Motion (SoM) (Wang et al., 2025a), struggle on extreme novel views far from the input trajectory, such as opposite-side views (view 1) or large angle offsets (view 2). We propose USplat4D, an Uncertainty-aware dynamic Gaussian Splatting model that produces more accurate and consistent 4D reconstruction. Left: rendered dynamic scene from our model for illustration alongside four sampled RGB inputs. Right: novel view synthesis at two extreme novel viewpoints. Please refer to supplementary video for clearer visual comparison.
@@ 1 Introduction
Reconstructing dynamic 3D scenes from monocular input is a fundamental problem across a variety of tasks, including augmented reality, robotics, and human motion analysis (Slavcheva et al., 2017; Li et al., 2023; Newcombe et al., 2015; Joo et al., 2014; Gao et al., 2022). However, despite its wide applicability, monocular dynamic reconstruction remains highly challenging (Liang et al., 2025c), particularly under occlusion and extreme viewpoint changes. Recently, The advent of 3D Gaussian Splatting (Kerbl et al., 2023) has enabled real-time photorealistic rendering and sparked a series of dynamic extensions (Luiten et al., 2024; Lei et al., 2025; Stearns et al., 2024; Liu et al., 2024; Duan et al., 2024; Huang et al., 2024; Yang et al., 2024c; b; Li et al., 2024; Sun et al., 2024b; Duisterhof et al., 2023; Das et al., 2024; Lin et al., 2024; Wu et al., 2025). These methods parameterize motion with shared canonical fields (Wu et al., 2024; Yang et al., 2024c; Liang et al., 2025b; Guo et al., 2024; Lu et al., 2024; Liu et al., 2024; Wan et al., 2024), deformation bases (Wang et al., 2025a; Das et al., 2024; Lin et al., 2024; Li et al., 2024), or direct 4D modeling (Duan et al., 2024; Yang et al., 2024b).
Despite their differences in formulation, existing dynamic Gaussian splatting methods often share a common assumption: motion is optimized uniformly across all Gaussians using 2D supervision such as depth (Yang et al., 2024a), optical flow (Teed & Deng, 2020), and photometric consistency (Doersch et al., 2023). This uniform treatment overlooks that some Gaussians are strongly constrained by recurring observations, while others are only weakly constrained. As a result, motion estimates drift under occlusion and synthesized views degrade at novel viewpoints.
@@ 2 Related Work
Dynamic Gaussian splatting. Recent advances in dynamic Gaussian Splatting have enabled monocular 4D reconstruction via per-Gaussian deformation or canonical motion modeling (Liang et al., 2025c; Yang et al., 2024c; b; Wu et al., 2024; Li et al., 2024) or high-fidelity dynamic scene reconstruction from multi-view inputs (Luiten et al., 2024; Wang et al., 2025b). To reconstruct 4D GS from the monocular video, methods such as SoM (Wang et al., 2025a), MoSca (Lei et al., 2025), Marbles (Stearns et al., 2024), and 4D-Rotor (Duan et al., 2024) use low-rank motion bases to regularize deformation, while others model canonical flows (Liang et al., 2025b; Liu et al., 2024). Although they demonstrate high-fidelity rendering on near-input validation views, they do not explicitly model the motion behind occluders or identify reliable Gaussians for motion guidance. MoSca (Lei et al., 2025) introduces a soft motion score but lacks structured propagation. In contrast, USplat4D selects high-confidence Gaussians and constructs an uncertainty-aware motion graph to propagate motion through spatio-temporally coherent connections, improving robustness in occluded regions and enabling localized refinement beyond low-rank modeling (Kim et al., 2024; Huang et al., 2024).
Uncertainty estimation in scene reconstruction. Uncertainty modeling has been widely explored in neural rendering (Li et al., 2022; Pan et al., 2022; Shen et al., 2021; 2022; Kim et al., 2022; Zhan et al., 2022; Yan et al., 2023; Lee et al., 2022; Sünderhauf et al., 2023), particularly to improve robustness under occlusion, sparse views, and ambiguity. To avoid overfitting on reconstructing the static scene, SE-GS (Zhao et al., 2025) designs an uncertainty-aware perturbing strategy by estimating the self-ensembling uncertainty. In dynamic settings, uncertainty has been used to smooth motion or reweight gradients (Kim et al., 2024), but typically as an auxiliary signal decoupled from the underlying motion representation. In contrast, USplat4D treats uncertainty as a central modeling component. We estimate confidence of each Gaussian and use it to guide key node selection, edge construction, and loss weighting in a spatio-temporal graph. This allows high-confidence Gaussians to guide motion propagation while reducing the influence of uncertain regions. To the best of our knowledge, this is among the first attempts and analyses to model the uncertainty and directly integrate it into graph-based motion modeling for dynamic reconstruction.
@@ 4.1 Dynamic Uncertainty Estimation
Existing dynamic Gaussian Splatting optimizes all primitives uniformly, even though some are well observed across time while others remain ambiguous. This causes drift under occlusion and instability at extreme viewpoints. We therefore assign each Gaussian a time-varying uncertainty to estimate its reliability per frame and use it to guide optimization in a model-agnostic way.
Per-Gaussian scalar uncertainty. A straightforward way to capture reliability is to assign each Gaussian $i$ a scalar uncertainty $u_{i,t}$ at every frame $t$. Intuitively, if a primitive is frequently and clearly observed, its uncertainty should be small; if it is rarely visible or weakly constrained, its uncertainty should be large. The photometric loss is defined as
@@
Here, $\Omega$ is the pixel index set, $\bar{C}_{t}^{h}$ and $C_{t}^{h}$ denote the ground-truth and rendered colors at pixel $h$, and ${c}_{i}$ is the color parameter of Gaussian $i$ at frame $t$. The rendered pixel color is obtained by $\alpha$-blending, where the blending weight is given by $T_{i,t}^{h}\alpha_{i}$, with $T_{i,t}^{h}$ the transmittance of Gaussian $i$ at pixel $h$, $\alpha_{i}$ its opacity, and $N_{g}$ is the number of Gaussians. By differentiating $\mathcal{L}_{2,t}$ with respect to ${c}_{i}$ and applying the local minimum assumption, we obtain the closed-form variance estimate (please see Appendix A.1 for detailed derivation):
@@
where $\Omega_{i,t}\subseteq\Omega_{t}$ denotes the set of pixels contributed by that Gaussian. We thus take this variance as the scalar uncertainty, i.e., $u_{i,t}:=\sigma_{i,t}^{2}$. However, the local minimum assumption may not hold everywhere. To account for unconverged pixels, we introduce an indicator function to test per-pixel convergence:
$\mathcal{I}(h)=\left\{\begin{array}[]{ll}1&\quad\textrm{if }\|\bar{C}_{t}^{h}-C_{t}^{h}\|_{1}<\eta_{c},\\ 0&\quad\textrm{otherwise},\end{array}\right$. (6)
where $\eta_{c}>0$ is a color-error threshold. For Gaussian $i$ at frame $t$, we define the aggregate indicator $\mathcal{I}_{i,t}=\prod_{h\in\Omega_{i,t}}\mathcal{I}_{t}(h),$ which equals $1$ only if all covered pixels are convergent. If $\mathcal{I}_{i,t}=0$, we assign a large constant $\phi$ to indicate high uncertainty. Therefore, the final scalar uncertainty is:
$u_{i,t}=\mathcal{I}_{i,t}\,\sigma_{i,t}^{2}+(1-\mathcal{I}_{i,t})\,\phi.$ (7)
This design jointly constraint strength with convergence status: Gaussians that are well supported by observations receive low $u_{i,t}$, reflecting high reliability, while unreliable ones are assigned high values, which enables trustworthy primitives to guide ambiguous ones during optimization.
@@ 4.2 Uncertainty-encoded Graph Construction
and connect $\mathcal{E}_{i}=\mathcal{E}_{j}\cup\{j\}$ and $j$ is the closet key nodes. Intuitively, each uncertain non-key node is attached to the most reliable key node that stays close to it over time, so its motion can be regularized by stable anchors for consistency. In both cases, uncertainty-aware kkNN ensures edges are anchored by reliable nodes, promoting stable motion propagation and preventing drift from uncertain regions.
@@ 4.3 Uncertainty-Aware Optimization
Vanilla optimization (Section 3) of dynamic Gaussians often fail under occlusion or extreme viewpoints. This leads unreliable primitives to drift, since they are optimized as strongly as reliable ones. To address this, we incorporate uncertainty into the objective: key nodes with stable observations serve as anchors, while non-key nodes are regularized more softly through interpolation. We design separate objectives for key and non-key nodes, then unify them in a total loss.
@@
$\mathcal{L}^{\textrm{non-key}}=\sum\nolimits_{t=0}^{T-1}\sum\nolimits_{i\in\mathcal{V}}\|\mathbf{p}_{i,t}-\mathbf{p}_{i,t}^{\mathrm{o}}\|_{\mathbf{U}_{w,t,i}^{-1}}+\sum\nolimits_{t=0}^{T-1}\sum\nolimits_{i\in\mathcal{V}}\|\mathbf{p}_{i,t}-\mathbf{p}_{i,t}^{\textrm{DQB}}\|_{\mathbf{U}_{w,t,i}^{-1}}+\mathcal{L}^{\textrm{motion,non-key}}.$ (13)
@@ 5 Experiments
5 scenes
1 $\times$ resolution
SC-GS (Huang et al., 2024) 14.13 0.477 0.49
@@ 1 $\times$ resolution
7 scenes
2 $\times$ resolution
Dynamic Gaussians (Luiten et al., 2024)   7.29 0.69
@@ 5.1 Datasets and Baselines
Baselines. We compare USplat4D with state-of-the-art dynamic Gaussian Splatting methods. We use SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025) as base models. SoM is widely adopted, and MoSca represents the current state of the art. Please see Appendix B for additional details.
@@ 5.2 Main Results
Results on DyCheck. Table 1 reports results on DyCheck validation views, which are close to the input trajectories. USplat4D consistently outperforms baselines across nearly all metrics; qualitative examples are shown in Figure 3. However, these validation views remain relatively easy. The more significant improvements emerge under extreme novel viewpoints (Figure 4), which are not included in Table 1 but are essential for assessing robustness under severe viewpoint shifts. Please see Appendix C.1 for extra baselines, and Appendix C.2 and C.3 for tracking and per-scene results.
@@
Results on Objaverse. Table 2 shows that USplat4D consistently surpasses SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025), with gains most pronounced under large viewpoint shifts ($120^{\circ}$$180^{\circ}$). Figure 6 confirms these improvements qualitatively: our method preserves geometry and textures under extreme novel views, where baselines often blur or collapse. We include more details in Appendix C.5, including dataset setups and per-scene results.
@@ 5.3 Ablation and Analysis
We conduct ablation studies to assess key design choices in USplat4D, using MoSca (Lei et al., 2025) as the base model and evaluating on the DyCheck validation set (Gao et al., 2022).
Table 3: Ablation study on uncertainty usage. We assess the impact of uncertainty estimation by removing it from key components in USplat4D individually.
@@
Table 3 shows that uncertainty is essential across key components of our framework. (a) When removed from key node selection and replaced with uniform 2D sampling, the graph fails to emphasize well-constrained Gaussians, leading to unstable anchors and degraded propagation. (b) Replacing UA-kkNN with distance-only kkNN weakens graph connectivity, as edges disregard node reliability and mistakenly connect unstable primitives. (c) Excluding uncertainty weighting from the training loss (i.e., applying the DQB loss without $\mathbf{U}$) reduces PSNR/SSIM, since unreliable Gaussians are updated as aggressively as reliable ones, causing drift across frames. Please see Appendix D additional discussion and other analysis.

### v2

@@ Abstract
Reconstructing dynamic 3D scenes from monocular input is fundamentally under-constrained, with ambiguities arising from occlusion and extreme novel views. While dynamic Gaussian Splatting offers an efficient representation, vanilla models optimize all Gaussian primitives uniformly, ignoring whether they are well or poorly observed. This limitation leads to motion drifts under occlusion and degraded synthesis when extrapolating to unseen views. We argue that uncertainty matters: Gaussians with recurring observations across views and time act as reliable anchors to guide motion, whereas those with limited visibility are treated as less reliable. To this end, we introduce USplat4D, a novel Uncertainty-aware dynamic Gaussian Splatting framework that propagates reliable motion cues to enhance 4D reconstruction. Our approach estimates time-varying per-Gaussian uncertainty and leverages it to construct a spatio-temporal graph for uncertainty-aware optimization. Experiments on diverse real and synthetic datasets show that explicitly modeling uncertainty consistently improves dynamic Gaussian Splatting models, yielding more stable geometry under occlusion and high-quality synthesis at extreme viewpoints. Project page: https://tamu-visual-ai.github.io/usplat4d/.
Figure 1: We show a challenging case from the DyCheck dataset (Gao et al., 2022), where a person casually rotates a backpack while being captured by a moving monocular camera. The goal is to reconstruct the dynamic object for arbitrary viewpoints and timestamps. The state-of-the-art dynamic Gaussian Splatting methods, e.g., Shape-of-Motion (SoM) (Wang et al., 2025a), struggle on extreme novel views far from the input trajectory, such as opposite-side views (view 1) or large angle offsets (view 2). We propose USplat4D, an Uncertainty-aware dynamic Gaussian Splatting model that produces more accurate and consistent 4D reconstruction. Left: rendered dynamic scene from our model for illustration alongside four sampled RGB inputs. Right: novel view synthesis at two extreme novel viewpoints. Please refer to the project page for clearer visual comparison.
@@ 1 Introduction
Reconstructing dynamic 3D scenes from monocular input is a fundamental problem across a variety of tasks, including augmented reality, robotics, and human motion analysis (Slavcheva et al., 2017; Li et al., 2023; Newcombe et al., 2015; Joo et al., 2014; Gao et al., 2022). However, despite its wide applicability, monocular dynamic reconstruction remains highly challenging (Liang et al., 2025c), particularly under occlusion and extreme viewpoint changes. Recently, the advent of 3D Gaussian Splatting (Kerbl et al., 2023) has enabled real-time photorealistic rendering and sparked a series of dynamic extensions (Luiten et al., 2024; Lei et al., 2025; Stearns et al., 2024; Liu et al., 2025; Duan et al., 2024; Huang et al., 2024; Yang et al., 2024b; c; Li et al., 2024; Sun et al., 2024b; Duisterhof et al., 2023; Das et al., 2024; Lin et al., 2024; Wu et al., 2025). These methods parameterize motion with shared canonical fields (Wu et al., 2024; Yang et al., 2024c; Liang et al., 2025b; Guo et al., 2024; Lu et al., 2024; Wan et al., 2024; Liu et al., 2025), deformation bases (Wang et al., 2025a; Das et al., 2024; Lin et al., 2024; Li et al., 2024), or direct 4D modeling (Duan et al., 2024; Yang et al., 2024b).
Despite their differences in formulation, existing dynamic Gaussian splatting methods often share a common assumption: motion is optimized uniformly across all Gaussians using 2D supervision such as depth (Yang et al., 2024a), optical flow (Teed and Deng, 2020), and photometric consistency (Doersch et al., 2023). This uniform treatment overlooks that some Gaussians are strongly constrained by recurring observations, while others are only weakly constrained. As a result, motion estimates drift under occlusion and synthesized views degrade at novel viewpoints.
@@ 2 Related Work
Dynamic Gaussian splatting. Recent advances in dynamic Gaussian Splatting have enabled monocular 4D reconstruction via per-Gaussian deformation or canonical motion modeling (Liang et al., 2025c; Yang et al., 2024b; c; Wu et al., 2024; Li et al., 2024) or high-fidelity dynamic scene reconstruction from multi-view inputs (Luiten et al., 2024; Wang et al., 2025b). To reconstruct 4D GS from the monocular video, methods such as SoM (Wang et al., 2025a), MoSca (Lei et al., 2025), Marbles (Stearns et al., 2024), and 4D-Rotor (Duan et al., 2024) use low-rank motion bases to regularize deformation, while others model canonical flows (Liang et al., 2025b; Liu et al., 2025). Although they demonstrate high-fidelity rendering on near-input validation views, they do not explicitly model the motion behind occluders or identify reliable Gaussians for motion guidance. MoSca (Lei et al., 2025) introduces a soft motion score but lacks structured propagation. In contrast, USplat4D selects high-confidence Gaussians and constructs an uncertainty-aware motion graph to propagate motion through spatio-temporally coherent connections, improving robustness in occluded regions and enabling localized refinement beyond low-rank modeling (Kim et al., 2024; Huang et al., 2024).
Uncertainty estimation in scene reconstruction. Uncertainty modeling has been widely explored in neural rendering (Shen et al., 2021; Li et al., 2022; Pan et al., 2022; Shen et al., 2022; Kim et al., 2022; Zhan et al., 2022; Yan et al., 2023; Lee et al., 2022; Sünderhauf et al., 2023), particularly to improve robustness under occlusion, sparse views, and ambiguity. To avoid overfitting on reconstructing the static scene, SE-GS (Zhao et al., 2025) designs an uncertainty-aware perturbing strategy by estimating the self-ensembling uncertainty. In dynamic settings, uncertainty has been used to smooth motion or reweight gradients (Kim et al., 2024), but typically as an auxiliary signal decoupled from the underlying motion representation. In contrast, USplat4D treats uncertainty as a central modeling component. We estimate confidence of each Gaussian and use it to guide key node selection, edge construction, and loss weighting in a spatio-temporal graph. This allows high-confidence Gaussians to guide motion propagation while reducing the influence of uncertain regions. To the best of our knowledge, this is among the first attempts and analyses to model the uncertainty and directly integrate it into graph-based motion modeling for dynamic reconstruction.
@@ 4.1 Dynamic Uncertainty Estimation
Vanilla dynamic Gaussian Splatting optimizes all primitives uniformly, even though some are well observed across time while others remain ambiguous. This causes drift under occlusion and instability at extreme viewpoints. We therefore assign each Gaussian $\mathbf{G}_{i}$ a time-varying uncertainty to estimate its reliability per frame and use it to guide optimization in a model-agnostic way.
Per-Gaussian scalar uncertainty. A straightforward way to capture reliability is to assign each Gaussian $i$ a scalar uncertainty $u_{i,t}$ at every frame $t$. Intuitively, if a primitive is frequently and clearly observed, its uncertainty should be small; if it is rarely visible or weakly constrained, its uncertainty should be large. The photometric loss is defined as:
@@
Here, $\Omega$ is the pixel index set, $\bar{C}_{t}^{h}$ and $C_{t}^{h}$ denote the ground-truth and rendered colors of pixel $h$ at frame $t$, and ${c}_{i}$ is the color parameter of Gaussian $i$. The rendered pixel color is obtained by $\alpha$-blending, where the blending weight is given by $T_{i,t}^{h}\alpha_{i}$, with $T_{i,t}^{h}$ the transmittance of Gaussian $i$ at pixel $h$, $\alpha_{i}$ its opacity, and $N_{g}$ is the number of Gaussians. By differentiating $\mathcal{L}_{2,t}$ with respect to ${c}_{i}$ and applying the local minimum assumption, we obtain the closed-form variance estimate (please see Appendix A.1 for detailed derivation):
@@
where $\Omega_{i,t}\subseteq\Omega_{t}$ is the set of pixels contributed to by that Gaussian. We thus take this variance as the scalar uncertainty, i.e., $u_{i,t}:=\sigma_{i,t}^{2}$. However, the local minimum assumption may not hold everywhere. To account for unconverged pixels, we introduce an indicator function to test per-pixel convergence:
$\mathds{1}_{t}(h)=\left\{\begin{array}[]{ll}1&\quad\textrm{if }\|\bar{C}_{t}^{h}-C_{t}^{h}\|_{1}<\eta_{c},\\ 0&\quad\textrm{otherwise},\end{array}\right$. (6)
where $\eta_{c}>0$ is a color-error threshold. For Gaussian $i$ at frame $t$, we define the aggregate indicator $\mathds{1}_{i,t}=\prod_{h\in\Omega_{i,t}}\mathds{1}_{t}(h),$ which equals $1$ only if all covered pixels are convergent. If $\mathds{1}_{i,t}=0$, we assign a large constant $\phi$ to indicate high uncertainty. Therefore, the final scalar uncertainty is:
$u_{i,t}=\mathds{1}_{i,t}\,\sigma_{i,t}^{2}+(1-\mathds{1}_{i,t})\,\phi.$ (7)
This design jointly adjusts the strength according to convergence status: Gaussians that are well supported by observations receive low $u_{i,t}$, reflecting high reliability, while unreliable ones are assigned high values, which enables trustworthy primitives to guide ambiguous ones during optimization.
@@ 4.2 Uncertainty-encoded Graph Construction
and connect $\mathcal{E}_{i}=\mathcal{E}_{j}\cup\{j\}$, where $j$ is the closest key nodes. Intuitively, each uncertain non-key node is attached to the most reliable key node that stays close to it over time, so its motion can be regularized by stable anchors for consistency. In both cases, uncertainty-aware kkNN ensures edges are anchored by reliable nodes, promoting stable motion propagation and preventing drift from uncertain regions.
@@ 4.3 Uncertainty-Aware Optimization
Vanilla optimization (Section 3) of dynamic Gaussians often fails under occlusion or extreme viewpoints. This leads unreliable primitives to drift, since they are optimized as strongly as reliable ones. To address this, we incorporate uncertainty into the objective: key nodes with stable observations serve as anchors, while non-key nodes are regularized more softly through interpolation. We design separate objectives for key and non-key nodes, then unify them in a total loss.
@@
$\mathcal{L}^{\textrm{non-key}}=\sum\nolimits_{t=0}^{T-1}\sum\nolimits_{i\in\mathcal{V}_{n}}\|\mathbf{p}_{i,t}-\mathbf{p}_{i,t}^{\mathrm{o}}\|_{\mathbf{U}_{w,i}^{-1}}+\sum\nolimits_{t=0}^{T-1}\sum\nolimits_{i\in\mathcal{V}_{n}}\|\mathbf{p}_{i,t}-\mathbf{p}_{i,t}^{\textrm{DQB}}\|_{\mathbf{U}_{w,i}^{-1}}+\mathcal{L}^{\textrm{motion,non-key}}.$ (13)
@@ 5 Experiments
5 scenes 1 $\times$ resolution SC-GS (Huang et al., 2024) 14.13 0.477 0.49
@@
7 scenes 2 $\times$ resolution Dynamic Gaussians (Luiten et al., 2024)   7.29 0.69
@@ 5.1 Setup
Baselines. We evaluate USplat4D against state-of-the-art dynamic Gaussian Splatting methods. We use SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025) as our main base models. Specifically, SoM is a widely adopted standard in the field, and MoSca represents the current state of the art. Notably, our framework is agnostic to the underlying architecture and is compatible with any method that estimates per-Gaussian motion. Please see Appendix B for additional implementation details (e.g., hyper-parameters for model training) and details of additional baseline methods.
@@ 5.2 Main Results
Results on DyCheck. Table 1 reports results on DyCheck validation views, which are close to the input trajectories. USplat4D consistently outperforms baselines across all metrics; corresponding qualitative examples are shown in Figure 3. However, these validation views remain relatively easy due to the proximity to the input views. The more significant improvements emerge under extreme novel viewpoints (Figure 4), which are not included in Table 1 but are essential for assessing robustness under severe viewpoint shifts. Our results avoids the collapse and distortions observed in the baselines. Please refer to Appendix C for the tracking results and per-scene comparison.
@@
Results on Objaverse. Table 2 shows that USplat4D consistently surpasses SoM (Wang et al., 2025a) and MoSca (Lei et al., 2025), with gains most pronounced under large viewpoint shifts ($120^{\circ}$$180^{\circ}$). Figure 6 confirms these improvements qualitatively: our method preserves geometry and textures under extreme novel views, where baselines often blur or collapse. Please refer to Appendix C for more quantitative and qualitative results.
Other results. We provide further results and analysis in Appendix C, including per-scene results on the DyCheck and Objaverse, evaluations on the NVIDIA (Yoon et al., 2020) and HyperNeRF (Park et al., 2021) datasets, results of the tracking task, and comparisons with additional baseline methods.
@@ 5.3 Ablation and Analysis
We conduct ablation studies to assess key design choices in USplat4D, using MoSca (Lei et al., 2025) as the base model and evaluating on the DyCheck validation set (Gao et al., 2022). As described in Section 4, our core design include the graph initialization strategy and the use of uncertainty cues during graph construction and optimization. Finally, we analyze the model’s capabilities.
Table 3: Ablation study on uncertainty usage and key node selection. We assess the impact of uncertainty estimation and key node selection strategy by removing them from key components in USplat4D individually.
@@
(d) w/o 3D gridization 19.50 0.712 0.25
Ablation study on uncertainty usage and key node selection. Table 3 shows that uncertainty and key node selection are essential across key components of our framework. (a) When removed from key node selection and replaced with uniform 2D sampling, the graph fails to emphasize well-constrained Gaussians, leading to unstable anchors and degraded propagation. (b) Replacing UA-kkNN with distance-only kkNN weakens graph connectivity, as edges disregard node reliability and mistakenly connect unstable primitives. (c) Excluding uncertainty weighting from the training loss (i.e., applying the DQB loss without $\mathbf{U}$) reduces PSNR/SSIM, since unreliable Gaussians are updated as aggressively as reliable ones, causing drift across frames. (d) Replacing the key node candidate sampling strategy from 3D grid-based sampling with spatially random sampling leads to performance degradation, since the non-uniform spatial distribution neither guarantees sufficient spatial coverage nor avoids redundant samples. Please see Appendix D for additional ablation study (on hyperparameter key / non-key ratio, color threshold, significant period threshold, and scaling factor) and discussion on time complexity.
Figure 7: Key node weight matrix and Gaussian segmentation. The weight matrix visualizes edge connections among key nodes. Nodes belonging to different diagonal blocks do not share weights, indicating the no connections across those blocks. The images further illustrate the corresponding Gaussian segmentation, where colors represent the block assignment of each key and non-key node.
Further analysis. Our model naturally supports multiple objects and inherently group Gaussians by motion. In key and non-key graphs, nodes with similar instance-level motion are strongly connected, while nodes from different instances have weak (or few) connections. As shown in Figure 7, the key-graph weight matrix (defined in Appendix A.2.1) becomes approximately block-diagonal after reordering nodes with standard community detection (e.g., spectral clustering (Von Luxburg, 2007)). After clustering key nodes, we assign non-key nodes to the same motion groups using their non-key graph connections. Each Gaussian gets a label, and when objects move differently, the segmentation closely matches instance segmentation with dynamic tracking. Besides, we further discuss the challenging cases including textureless region, fast motion, and deforming objects in Appendix E.